Name of Project

Development of a Public Domain High Performance Fortran Translation System.

Lead Site

The University of Southampton

Other Consortia and Partnership Members

JISC Funding Sought in Year 1

25K

Total Program Funding Sought

Outline Description

High Performance Fortran (HPF) is the new language standard for developing Scientific programs to run on distributed memory supercomputers. It consists of a set of extensions to Fortran 90. Whilst several software vendors have announced implementations of HPF, with promised availability in the near future, we propose to develop a portable HPF translation system which would be made freely available to HE establishments.

Importing expertise from the University’s Department of Electronics and Computer Science, where research on HPF and precursor systems has been pursued for several years, we can credibly produce a prototype system within the academic year 93/94. Follow-on funding would allow the prototype to be developed to a product standard, and subsequently improved and maintained.